OpenStudy (anonymous):

the x intercept of the line is -5 and the y intercept of the line is -2 what is the equation of the line

ganeshie8 (ganeshie8):

equation of line wid x intercept = a, y intercept = b : \(\huge \frac{x}{a} + \frac{y}{b} = 1\)

ganeshie8 (ganeshie8):

simply plug the intercepts in above equation, and simplify

OpenStudy (anonymous):

so is the answer y=-5/2-5 y=2/5x+2 y=5/2x-2 y=-2/5x-2

OpenStudy (anonymous):

@ganeshie8